OpenStudy (anonymous):

guys, im really bad at math. will someone please explain rational expressions to me?!! i don't understand how to simplify them, well.. i pretty much dont understand anything about them. please help?

OpenStudy (anonymous):

http://www.purplemath.com/modules/rtnldefs2.htm

OpenStudy (anonymous):

an example would be best you got one?

OpenStudy (anonymous):

Simplify: \[\frac{9ab^8} {2a^9 b^11} \times \frac {a^8b^8}{3a^3b^3}\]

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

If I did this correctly, then your answer should look like this..

OpenStudy (anonymous):

Double checking for you right now. c:

OpenStudy (anonymous):

i am going to guess that it is \[\frac{9ab^8} {2a^9 b^{11}} \times \frac {a^8b^8}{3a^3b^3}\]

OpenStudy (anonymous):

there are many ways to do it, but perhaps the simplest is first to multiply by adding the exponents on the like terms, go ge t \[\frac{9a^9b^{16}}{6a^{12}b^{14}}\]

OpenStudy (anonymous):

now if your job is to write with positive exponents only go to \[\frac{3b^{16-14}}{2a^{12-9}}=\frac{3b^2}{2a^3}\]

OpenStudy (anonymous):

there are other ways to do it. for example you could have cancelled the 9 and the 3 at the first step

OpenStudy (anonymous):

My answer that was given before simplifies down to: \[\frac{ 3b ^{12} }{ 2a ^{3} }\]

OpenStudy (anonymous):

yeah i rewrote it because i think there was a typo in the question